### v1

Abstract
Quantum Hall effects provide intuitive ways of revealing the topology in crystals, i.e., each quantized “step” represents a distinct topological state. Here, we seek a counterpart for “visualizing” quantum geometry, which is a broader concept. We show how geometry emerges in quantum as an intrinsic consequence of unitary evolution, independent of specific details or approximations, suggesting quantum geometry may have widespread applicability. Indeed, we exemplify geometric observables, such as oscillation, dephasing, in spin and band scenarios. These phenomena are robust owing to the continuity of geometry, and can be tuned by geometric parameters. Anomalies, supported by both analytic and numerical solutions, underscore the advantages of adopting a geometric perspective, potentially yielding distinguishable experimental signatures.
pacs:
Introduction. Quantum geometry, as a rising field 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 ; 11 , is aimed to reveal geometry’s importance in quantum. This geometry is irrelevant to the shapes of materials or other tangible geometries seen in our real space, rather it is related to the structure of quantum states 12 , or in limited senses, related to geometric tensor $T_{i,j}$ defined in $k$-space 1 ; 3 ; 4 ; 5 . A hallmark for such abstract geometry is the quantum Hall state, indexed by Chern numbers, 11a ; 11b which characterize a series of discrete states, namely topological states. A material’s topology can be “visualized” by Hall conductivities: each quantized step of ${\sigma}_{xy}$ corresponds to a distinct topology. This quantization displays a universal $e^{2}/h$, independent of material species.
Compared with topology, quantum geometry is an even broader concept 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 24 ; 25 ; 26 ; 27 , referring to generic properties derived from surfaces, curves, etc., either locally or globally, whether quantized or not. For example, the quantum metric 7 ; 8 ; 13 ; 14 is recently studied for optical transitions 6 ; 7 ; 15 and flat-band superconductivity 16 ; 17 , with goals from understanding strange metallicity, 19 correlated phases, 14 ; 20 ; 21 ; 22 to unifying seemingly unrelated domains like heavy Fermion 10 ; 23 and Moiré super-lattices 19 ; 24 ; 25 . Essentially, these questions are posed in analog with topology: does quantum geometry cause any robust phenomenon? How to visualize the geometry to show its existence?
In this work, we examine an unremarked consequence of unitarity of evolution and show that geometry directly arises from Schrödinger equation (SE). Robust features such as oscillation, dephasing (in spin and band scenarios) make quantum geometry visible, just as the quantized ${\sigma}_{xy}$ does for topology. Some features bring insights to the conventional belief, just like the ${\sigma}_{xy}$ “steps” motivate us to think more about the classical picture.
The foundation of geometry. Geometric notions are introduced based on a myriad of presumptions: adiabaticity 1 , semi-classical wave packages 3 ; 18 , thermodynamic potentials 8 ; 16 ; 17 . Although there are certainly pragmatic values, it raises concerns: (1) for a particular notion, its validity is limited by these preconditions; (2) in view of diverse proposals, one may wonder if the geometry is a convenient “metaphor” rather than a fundamental concept. Here, we will introduce geometry based on the mere fact that the evolution operator is unitary, free of further assumptions.
Unitary evolution is usually emphasized for its preservation of a state vector’s magnitude, i.e., probability, while its geometric meanings as rotations are often overlooked. For example, evolution operator $U$ for spin 1/2 (2D Hilbert space) belongs to SU(2) that involves three Euler angles ${\phi},{\theta},{\psi}{\in}[0,2{\pi})$. That means in addition to the dynamic parameters (time $t$ and parameters contained in Hamiltonian $H$), $U$ could be labeled by angles, such as $U({\phi},{\theta},{\psi})$. By setting $U(t;B,{\omega},{\cdots})=U({\phi},{\theta},{\psi})$, we get smooth maps between the two types of parameters: $(t;B,{\omega},{\cdots}){\mapsto}({\phi},{\theta},{\psi})$, called $Z$-map denoted as
$\begin{split}Z:X_{e}{\rightarrow}X_{g}~{}.\end{split}$ (1)
$X_{e}$ and $X_{g}$ stand for the spaces of dynamic and geometric parameters. Interestingly, there is intrinsic mismatch between them: $X_{e}$ often contains boundless parameters, such as $t,~{}B{\in}(-{\infty},+{\infty})$, while angles in $X_{g}$ are bounded and periodic. Intuitively, the $Z$-map is about linking two objects of different “shapes”, like a large piece of paper to wrap up an orange. Accurately, a $Z$-map often connects a non-compact space to a compact one, characterized by the dimensions and natures of $X_{e}$ and $X_{g}$. Restricting to $\text{dim}X_{e}=\text{dim}X_{g}$, we find 1D is simple, just winding around (mod $2{\pi}$, Fig. 1a). In 2D, three topologies of $X_{e}$ are listed in Fig. 1b-d; in some of them, singularity might be inevitable, such as mapping from cylinder (Fig. 1c) or torus (Fig. 1d) to $S^{2}$ (a sphere surface), and these singularities may have important indications.
Evidently, the single SE is made into numerous categories based on the shapes of $X_{e}$ and $X_{g}$ — geometry and topology have sneaked in for the mere sake of $U$ being unitary. Under $Z$-map, for instance, by setting Brillouin zone (BZ) to be $X_{e}$ (Fig. 1d), one is led to find various geometric quantities defined in $k$-space such as Berry curvature; one may also find the existence of Berry phase closely related to the singular pole in Fig. 1c. Here, however, we focus on a simple isomorphic case Fig. 1b (free of singularity) to demonstrate robust abnormality that merely relies on continuity in geometry, without referring to specific geometric quantities.
Figure 1: (color online): $Z$-map (restricted to $X_{g}{\cong}S^{2}$) under different parameterizations of Hamiltonian. 1D (a) and 2D: (b) isomorphism $S^{2}{\rightarrow}S^{2}$; (c) plane $\mathbb{R}^{2}{\rightarrow}S^{2}$; (d) Torus $T^{2}{\rightarrow}S^{2}$. Singular points are inevitable for (c)(d) for distinct topologies. For (c), imagine a big piece of paper wrapping (winding around) an orange and encountering an annoying point at the pedicle. For (d), surgery has to be performed on the sphere to make it become a torus.
Consider a $Z$-map based on a spin driven by “pie slice” loops of $\textbf{B}(t)$ (Fig. 2a), with two straight and one arc sections. Bearing in mind that the goal is 2D $Z$-map like Fig. 1b, we recognize $X_{e}:=({\Theta},{\omega})$, an unbounded ${\omega}:=2{\pi}/{\tau}$ ($\tau$ is the cycle period) and a bounded $\Theta$ (the vertex angle); other parameters are fixed. Such a choice makes $X_{e}{\cong}S^{2}$, where $\cong$ means isomorphic.
Importantly, under $Z$-map, a spirit is expressing one-cycle evolution $\mathcal{U}$ in geometric parameters in $X_{g}$ instead of the original $X_{e}$, which corresponds to re-interpretation of $\mathcal{U}$ as rotations. Technically, $\mathcal{U}=\mathcal{R}{\cdot}{\Lambda}$. (Supporting information, SI), where rotation $\mathcal{R}$ ensures the resulting spin will align with the desired ray (orientation), and $\Lambda$ is a diagonal matrix to adjust the phase. It can be shown (SI) spin 1/2 will lead to
$\begin{split}\mathcal{U}({\Theta},{\Phi})=\begin{pmatrix}\text{cos}(\frac{\Theta}{2})e^{-i{\Phi}}&-\text{sin}(\frac{\Theta}{2})e^{i{\Phi}}\\ \text{sin}(\frac{\Theta}{2})e^{-i{\Phi}}&\text{cos}(\frac{\Theta}{2})e^{i\Phi}\end{pmatrix}\end{split}$ (2)
For spin 1 (a 3D Hilbert space), $\mathcal{U}({\Theta},{\Phi})$ is equal to
$\begin{split}\begin{pmatrix}\frac{1}{2}e^{-i{\Phi}}(1+\text{cos}({\Theta})&-\frac{\sqrt{2}}{2}\text{sin}({\Theta})&\frac{1}{2}e^{i{\Phi}}(1-\text{cos}({\Theta}))\\ \frac{\sqrt{2}}{2}e^{-i{\Phi}}\text{sin}({\Theta})&\text{cos}({\Theta})&-\frac{\sqrt{2}}{2}e^{i{\Phi}}\text{sin}({\Phi})\\ \frac{1}{2}e^{-i{\Phi}}(1-\text{cos}(\Theta))&\frac{\sqrt{2}}{2}\text{sin}({\Theta})&\frac{1}{2}e^{i{\Phi}}(1+\text{cos}({\Theta}))\end{pmatrix}\end{split}$ (3)
Equations.  2, 3 are based on $X_{g}:=({\Theta},{\Phi}){\cong}S^{2}$. Given the remaining parameters are fixed, a $Z$-map (just like Fig.  1b) is built
$\begin{split}Z:({\Theta},{\omega}){\mapsto}({\Theta},{\Phi}).\end{split}$ (4)
Both $\Theta$ and $\Phi$ are geometric parameters. The $\Theta$ refers to the vertex angle of $\textbf{B}(t)$, while $\Phi$ is a quantum phase angle arising from evolution (SI). Thus, the dynamic information in Eq. 2, 3 is solely in $\Phi$. Our goal is to show observables are only dependent on $\Theta$. Then, the difficult solving for ${\Phi}(t;B,{\omega}{\cdots})$ is eluded 27a ; more importantly a geometric effect is revealed.
Quantum oscillation & dephasing for spins. Consider a magnetic resonance experiment: a spin in a cyclic field $\textbf{B}(t)$ (Fig. 2a). In reality, if there is deviation from the model, e.g., “twists” happen in the straight section, or “bumps” exist in the arc section (Fig. 2b). We can show the forms of $\mathcal{U}$ in Eq. 2, 3 maintain but only ${\Theta}$ or $\Phi$ take slightly different values. (SI) Thus, the present analysis is insensitive to details of $\textbf{B}(t)$ functions.
The observable is spin population over eigenstates $|s,s_{z}{\rangle}$ after $t{\to}{\infty}$ (cycle number $n{\to}{\infty}$). Given an initial state is spin up, i.e., ${\varphi}(t=0)=|\frac{1}{2},\frac{1}{2}{\rangle}$ or $|1,1{\rangle}$. The average population of $n$ cycles is
$\begin{split}p_{n}(s,s_{z}):=\frac{1}{n}{\sum}_{j}^{n}|{\langle}s,s_{z}|\mathcal{U}^{j}|{\varphi}(0){\rangle}|^{2}.\end{split}$ (5)
For example, $p_{n}(1,-1)$ means the pumping probability of $|s=1,s_{z}=1{\rangle}{\to}|s=1,s_{z}=-1{\rangle}$. Using Eq. 2, 3, we (numerically) evaluate $p_{G}:=p_{n{\to}{\infty}}{\approx}p_{n=100}$ against $({\Theta},{\Phi})$, presented in the right half of Fig. 2c, d. (The phase diagram is symmetric with ${\Phi}=0$). Since $\Phi$ is periodic, $({\Theta},{\Phi})$ actually forms $S^{2}$ surface. If the system traverses along latitudes of $({\Theta},{\Phi})$, it should periodically encounter maximum $p_{G}$, forming an oscillation.
One possible realizing is to control the cycle period ${\tau}$ (or frequency $f=1/{\tau}$). Roughly, the shift of quantum phase ${\Delta}{\Phi}=E{\cdot}{\Delta}{\tau}/{\hbar}$. The energy $E{\sim}{\mu}_{B}\bar{B}$, where $\bar{B}$ is the average field in one cycle. Supposing $\bar{B}=0.01$ T, to observe the oscillation, we need frequency in MHz ranges, which is quite achievable with current techniques. 28 ; 29 ; 30 ; 31 ; 32 ${\Delta}{\Phi}=2{\pi}{\cdot}1.4{\times}10^{8}{\cdot}(1/f-{\tau}_{0})$, where suppose ${\tau}_{0}{\sim}10^{-7}$ s ($f_{0}{\sim}10$ MHz) is a time period making ${\Phi}=0$. Since $\Phi$ is periodic, there are infinite many ${\tau}_{0}$. Here we just choose the MHz range to make a plot. Note that the adiabaticity in the arc section requires $hf_{0}{\ll}\bar{\Delta}$. The current parameters lead to $hf_{0}{\sim}10^{-27}~{}{\ll}~{}\bar{\Delta}{\sim}10^{-25}$ J, satisfying the precondition.
Observations. (1) Pumping oscillates with $\textbf{B}(t)$’s frequency (Fig. 2e, f), which has a geometric origin: the system hovers around the latitudes of $S^{2}$ (inset of Fig. 2e). (2) Spin 1/2 exhibits an oscillation frequency double with spin 1 under a common condition (Fig.2e, f), which is related to that half-integer $s$ gives -1 after rotating $2{\pi}$, while integer $s$ gives +1. (Sec. 2.2 of SI). (3) The oscillation is switched on/off (Fig. 2e, f) by tuning the geometric angle of the $\textbf{B}(t)$ loop (Fig. 2a). A flat line (as ${\Theta}{\to}{\pi}$) means pumping probability is insensitive to the quantum phase ${\Phi}$. A basic feature in quantum is the phase’s influence (e.g., quantum interference); now the phase becomes effectless, as though it disappears. The dephasing is shared feature for half-integer and integer spins (Fig. 2d, e), and it only depends on ${\Theta}$, thus we call it geometric dephasing.
Figure 2: (color online): (a) A standard loop and its distortion due to deviation from the ideal. (b) A quenched pumping ($p_{G}{\to}0$) is geometrically interpreted as approaching to N-pole of $S^{2}$, and dephasing is approaching the S-pole. (c)(d) Comparison of analytic (left) and numerical solutions (right, at each pixel $p_{G}{\approx}p_{n=100}$). (e)(f) Quantum oscillation ($s=1/2,1$) approaches to dephasing tuned by angle ${\Theta}:0{\to}{\pi}$.
$Z$-map not only helps demonstrate quantum oscillation and dephasing, but also provides a pathway of finding analytic solutions for $p_{G}$ (Sec. 2 of SI). The idea is mapping the evolution path to a special subspace of $X_{g}$, namely ergodic subgroup, 26 ; 27 in which probability density is uniform according to quantum Liouville’s theorem recently proved 27 . With integration approaches, we obtain
$\begin{split}p_{G}(1,-1)=\frac{3}{8}\frac{\text{sin}^{4}(\frac{\Theta}{2})}{(1-\text{cos}^{2}(\frac{\Theta}{2})\text{cos}^{2}(\frac{\Phi}{2}))^{2}}\end{split}$ (6)
$\begin{split}p_{G}(1,0)=\frac{\text{sin}^{2}(\frac{\Theta}{2})}{1-\text{cos}^{2}(\frac{\Theta}{2})\text{cos}^{2}(\frac{\Phi}{2})}-\frac{3}{4}\frac{\text{sin}^{4}(\frac{\Theta}{2})}{(1-\text{cos}^{2}(\frac{\Theta}{2})\text{cos}^{2}(\frac{\Phi}{2}))^{2}}\end{split}$ (7)
Analytic and numerical $p_{G}$ are compared in Fig. 2c, d. The analytic solutions confirm the dephasing, as Eq. 6, 7 become independent of $\Phi$ at ${\Theta}={\pi}$. It also proves the exact limit values: $p_{G}(1,-1){\to}\frac{3}{8}$ and $p_{G}(1,0){\to}\frac{1}{4}$ at $\Theta{\to}{\pi}$.
The pumping $p_{G}{\neq}0$ observed here is based on ${\hbar}{\omega}{\ll}\bar{\Delta}$ except for an infinitesimal period of gap closing in the cycle, making it different from the adiabatic evolution where ${\hbar}{\omega}{\ll}\bar{\Delta}$ constantly holds, also different from Rabi oscillations, 33 which lead to pumping in the regime of ${\hbar}{\omega}{\sim}\bar{\Delta}$.
Quantum oscillations for bands. Given $k$ is conserved, similar analysis can be made on $H(k)$ as a (continuous) series of spin models labelled by $k$. Phonon plays the roles of cyclic fields, and $\textbf{B}=0$ corresponds to gap closing, which might lead to band inversion and thus topological phase transition (TPT). The spin observable is replaced with inter-band pumping, a charge distribution near gap closing $k_{0}$, formed after a number of TPTs driven by phonons. 29 ; 30 ; 31 ; 32 Consider a 1D two-band model (Fig. 3a)
$\begin{split}H(k)=c_{H}\begin{pmatrix}-{\epsilon}(t)-\text{cos}(k)&-i\text{sin}(k)\\ i\text{sin}(k)&{\epsilon}(t)+\text{cos}(k)\end{pmatrix},\end{split}$ (8)
where ${\epsilon}(t)={\epsilon}_{0}+A_{ph}{\cdot}\text{sin}({\omega}t)$ depicts phonon’s effects.
$Z$-map for bands relies on numerical methods to determine (${\Theta},{\Phi}$ are now functions of $k$). Given the model Eq. 8, we employ Trotter decomposition 34 ; 35 to evaluate $U$. Basically, the method discretizes the evolution into steps. Once $U$ is obtained, we use Eq. 2 to deduce $({\Theta},{\Phi})$. The $Z$-map is graphed in Fig. 3b, i.e., $H(k)$ is mapped to $S^{2}$ surface (inset of Fig. 3b).
Now $k$ plays roles as ${\omega}$ in spins, tracing out a path on $S^{2}$. Analytic solutions (Fig. 3c) suggest the maximum $p_{G}$ occurs at ${\Phi}=0$. Thus, when the longitude ${\Phi}=2n{\pi}$ (dashed lines in Fig. 3b) is passed, a peaked probability density occurs. However, as ${\Theta}{\to}0$, the peak width vanishes. Thus, only the first few orders have enough chance to be detected. The numerical simulations are presented in Fig. 3d. In principle, the peaks are equal heights, but for finite $k$ resolution ($k$-sampling eludes the peak), they appear decreasing. Physically, that means the probability density does not decay, while probability does.
In analog to spin pumping oscillating with $\omega$, band pumping should oscillate in $k$-space (Fig. 3d), for the same reason of hovering around $S^{2}$ surface. The oscillation is not limited by specific phonons, as long as gap is closed and TPT will happen. Indeed, different phonon periods ${\tau}_{ph}$ (representing dynamics of different modes) exhibit similar patterns, except for the longer ${\tau}_{ph}=3$ ps leading to a larger quantum phase $\Phi$ and thus passing more rounds across $S^{2}$ than ${\tau}_{ph}=1$ ps (Fig. 3d).
On one hand, the oscillating is contrary to results from perturbation, such as Fermi golden rule (FGR), which suggests pumping ${\propto}|\hat{V}|^{2}{\delta}({\omega}-\bar{\Delta}/{\hbar})$. The $\delta$-function is usually treated as a Lorentz function, indicating that ${\delta}({\omega}-\bar{\Delta}/{\hbar})$ monotonously decrease as $\bar{\Delta}$ increases. Thus, when $\hbar{\omega}$ is apparently lower than $\bar{\Delta}$ (here, ${\hbar}{\omega}{\sim}4$ meV ${\ll}\bar{\Delta}$), pumping monotonously decays as $k$ moves away from $\Gamma$, i.e., pumping has a maximum at $\Gamma$, where $\bar{\Delta}(k=0)$ has a minimum. However, $Z$-map suggests (1) an oscillating picture rather than monotonous trend (2) the maximum is not at $\Gamma$ (in fact, $\Gamma$ gives a local minimum of $p_{G}$). Previous simulations 34 caught such features but did not raise much attention. Moreover, the pumping proves fractional ($p_{G}{\leq}1/2$), a feature absent in perturbation.
On the other hand, with a coarse view (average to smooth out $p_{G}(k)$), the envelop of the oscillations will reduce to a shape like a Lorentz function centered at the gap closing point. Nonetheless the differences between $Z$-map and perturbation suggest (a) the higher orders discarded by perturbation might provide additional features; (b) the gap closing (causing TPT), even it is super shot and hardly affects the average gap, might provide a non-trivial influence.
Geometry provides a re-interpretation of phenomena. Conventionally, we understand a vanishing pumping from an energetic point view: ${\hbar}{\omega}{\ll}\bar{\Delta}$ making ${\delta}({\omega}-\bar{\Delta}/{\hbar}){\to}0$. In this limit, band details would not matter. $Z$-map gives a geometric understanding: ${\hbar}{\omega}{\ll}\bar{\Delta}$ is an approaching to the N-pole of the $S^{2}$, and since energetics correspond to motion along latitudes, it shows no motion near the N-pole, thus causing no effect. This is similar to the S-pole situation, which leads to dephasing ($p_{G}{\neq}0$). Thus, quantum geometry unifies seemingly unrelated phenomena: dephasing and disappearance of pumping are both interpreted as approaching to the poles of $S^{2}$.
Since detecting pumping with $k$-resolution could be stringent, to facilitate experimental observation, we make a proposal only dependent on the total probability (sum over BZ). The idea is to employ local site energy ${\epsilon}_{0}$ (instead of $k$) to drive the oscillation. A plausible tuning for ${\epsilon}_{0}$ is bias voltage or strain; a suitable system could be ZrTe5 29 ; 30 a narrow-gapped topological insulator at the vicinity of TPT. The only drawback is averaging might reduce the contrast; indeed, Fig. 3e shows the pumping probability is 1%. Fortunately, the pump-probe techniques 29 ; 32 provide sensitive detection of charge pumping by measuring the change of reflectivity ${\Delta}R$ with resolution ${\Delta}R/R$ $10^{-5}{\sim}10^{-6}$ n4 . Given ${\tau}_{ph}=1$ ps, and ${\epsilon}_{0}=-0.95$ eV, we calculate the total $p_{G}$ in Brillouin zone (Fig. 3e).
Figure 3: (color online): (a) Band structures of $H(k)$. After cyclic TPT, stable population is established near gap closing (near $\Gamma$, orange region). (b) The $k$ is interpreted as the coordinate of a continuous path in $({\Theta},{\Phi})$ space ($S^{2}$ surface shown by the inset). Every time longitude ${\Phi}=0$ (dashed) is passed, pumping probability $p_{G}$ gets to maximum (1st, 2nd, 3rd…). The color scales of dots stand for $k$’s value in unit of $\pi$. (c) Analytic solutions of $p_{G}$ (for spin 1/2). (d) Numerical results of $p_{G}$ in B.Z. ($p_{G}(k)=p_{G}(-k))$ subject to different phonon period ${\tau}_{ph}$, and its correspondence to $({\Theta},{\Phi})$ (lower panels). Maximum of $p_{G}$ occurs at dashed line (lower panels) crossing $\Phi$ (orange dots). Since $\Phi$ is periodic, the “jump” from -1 to 1 is actually continuous. (e) Oscillation of pumping tuned by bias voltages.
Conclusion. Our goals are twofold. First, elucidate the origin of geometry, endowing its entrance with a generic source, unconstrained by specific scenarios. In view of existing proposals involving different levels of quantum, statistics, etc., 2 ; 3 ; 4 ; 7 ; 8 ; 9 we are not expecting the framework is universal, but it might furnish a more unified one. Second, show the existence and influence of quantum geometry with observables in spin and band. Intuitively speaking, topology is demonstrated by “steps”, and we find geometry could be demonstrated by some “wavy lines”; moreover, these wavy lines can be tuned (for spin, by the loop vertex angle $\Theta$, and for band by band inversions, i.e., the presence or absence of TPT 26 ; 29 ), such that their origin could be tested and confirmed. The technique $Z$-map is useful for assigning a geometric interpretation, e.g., oscillation is circling around $S^{2}$ and dephasing is approaching to the poles. It is also useful in finding analytic solutions Eqns. 6,  7.